# Chen and Peng

### v2

@@ 5.4 Distance Weighted by Stationary Distribution
In other words, the state action pair $({\bm{s}},{\bm{a}})$ with higher occupancy measure ${\mu^{\pi}_{AB}}({\bm{s}},{\bm{a}})$ contributes more to the overall decomposition error. We note that the error bound in $\mu$-norm is weaker than that in $\ell_{\infty}$-norm. Nevertheless, a stationary distribution weighted error bound is sufficient in many practical scenarios. Similar ideas are also quite common in policy evaluation literature (Cai et al. 2019, Tsitsiklis and Van Roy 1996, Bhandari et al. 2021).
@@ 7 Applications of Markov Entanglement
This section applies Markov entanglement theory to practical scenarios. We first show how specific MDP structures simplify entanglement analysis and produce sharp decomposition error bounds. We then demonstrate how Markov entanglement provides an efficient test criteria of value decomposition for practitioners.
@@
Weakly-coupled MDPs (WCMDP) are a rich class of multi-agent model that capture many real-world applications such as supply chain management, queuing network and resource allocations (Adelman and Mersereau 2008, Brown and Zhang 2023, Shar and Jiang 2023). Compared to general multi-agent MDP, WCMDP further ensures each agent follow its local transition while the agents’ actions are coupled with each other. Formally,
@@
We demonstrate the proof for two-agent WCMDP and the generalization to multi-agent WCMDP is straightforward. Consider ${\bm{P}}_{A}^{\pi^{\prime}}$ be the transition of agent $A$ under local policy $\pi^{\prime}$. We focus on agent $A$
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{2}\sum_{{\bm{s}},{\bm{a}}}{\mu^{\pi}_{AB}}({\bm{s}},{\bm{a}})\sum_{s_{A}^{\prime},a_{A}^{\prime}}\left\lvert\sum_{s_{B}^{\prime}}P^{\pi}_{AB}({\bm{s}}^{\prime},a_{A}\mid{\bm{s}},{\bm{a}})-\sum_{s_{B}^{\prime}}P({\bm{s}}^{\prime}\mid{\bm{s}},{\bm{a}})\pi^{\prime}(a_{A}^{\prime}\mid s_{A}^{\prime})\right\rvert$
@@
=$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{2}\sum_{{\bm{s}}^{\prime}}{\mu^{\pi}_{AB}}({\bm{s}}^{\prime})\sum_{a_{A}^{\prime}}\left\lvert\pi(a_{A}^{\prime}\mid{\bm{s}}^{\prime})-\pi^{\prime}(a_{A}^{\prime}\mid s_{A}^{\prime})\right\rvert\,.$
@@ 7.1.1 Index Policies are Asymptotically Separable
We further dive into the more structured multi-agent MDPs and introduce the following model of Restless Multi-Armed Bandit (RMAB), a special instance of weakly coupled MDP which is also widely used in operations research literature (Whittle 1988, Weber and Weiss 1990, Gast et al. 2023, Zhang and Frazier 2021, 2022).
@@ A Restless Multi-Armed Bandit is an $N$-agent WCMDP that further satisfies
  • There are two available actions for each agent: 00 for idle and $1$ for activate.
@@
In other words, RMAB is WCMDP with two actions and homogeneous agents that are coupled under constraint $\sum_{i=1}^{N}a_{i}=M$ at any global state $s_{1:N}$. This coupling of agents is often referred to budget constraint. For example, healthcare platform can reach out only a fraction of patients at a time due to the cost budget of interventions (Baek et al. 2023).
@@
The index policies trace back to the well-known optimal Gittins Index (Weber 1992) and asymptotic optimal Whittle Index (Whittle 1988, Weber and Weiss 1990, Gast et al. 2023). More recent work generalizes Whittle Index to fluid based index policies (Verloop 2016, Gast et al. 2024). However, computing the best index policy typically requires the knowledge of system transition a prior. As an alternative, Qin et al. (2020), Azagirre et al. (2024), Baek et al. (2023), Nakhleh et al. (2021), Wang et al. (2023), Avrachenkov and Borkar (2022) apply data-driven method to optimize the index policy. Among these work, Qin et al. (2020), Azagirre et al. (2024), Baek et al. (2023) report great empirical success in industry-scale implementations. Understanding the mystery behind such success calls for a theory for general index policies. We then present our main result for this subsection
@@
Theorem 7.6 requires two standard technical conditions for index policies: non-degenerate and uniform global attractor property (UGAP), which restrict chaotic behavior in asymptotic regime and will be detailed in Appendix I. We note here these two assumptions are also used in many previous theoretical work on index policies (Weber and Weiss 1990, Verloop 2016, Gast et al. 2023, 2024).
@@
This sublinear error result explains why the value decomposition in Qin et al. (2020), Azagirre et al. (2024), Baek et al. (2023) manages to effectively approximate the global value function in large-scale practical applications. It also justifies Meta Algorithm 1 as an effective approach to evaluating or comparing index policies.
@@
This lemma builds upon Proposition 7.2. We thus focus on the deviation from ${\bm{m}}$ to ${\bm{m}}^{\ast}$. We extend the concentration analysis from Gast et al. (2023, 2024) to derive a new stability bound for the RHS. Specifically, we finishing the proof via demonstrating the deviation decays at the rate ${\mathcal{O}}(1/\sqrt{N})$.
@@ 7.2 Efficient Verification of Value Decomposition
For practitioners, verifying the feasibility of value decomposition remains a significant challenge. Typically, they have to rely on indirect methods—for instance, evaluating policies derived from value decomposition in real-world settings or simulation environments, as seen in Baek et al. (2023), Qin et al. (2020), Azagirre et al. (2024). While these policies often outperform baselines, it is unclear whether the learned decompositions accurately approximate the true global values. Moreover, real-world verifications are often prohibitively expensive. For example, the Lyft experiment (Azagirre et al. 2024) required a complex time-split design and posed substantial software engineering challenges in reliability and stability. Similarly, Baek et al. (2023), Qin et al. (2020) depended on extensive offline data collection to construct viable simulation environments.
@@
In other words, we can apply a Monte-Carlo estimation for estimating $E_{A}({\bm{P}}^{\pi}_{AB})$. Notice Eq. (9) is convex for ${\bm{P}}_{A}$ and this optimization can be solved distributively at each $(s_{A},a_{A})$ pair, which enables efficient solutions. Moreover, Eq. (9) only requires the knowledge of one-step transition ${\bm{P}}^{\pi}_{AB}(s_{A}^{\prime},a_{A}^{\prime}\mid{\bm{s}}^{t},{\bm{a}}^{t})$, which can often be easily calculated or simulated. For (weakly-)coupled MDPs, we can apply Proposition 7.2 and further eliminate the transition term,
@@
7.2.1 Numerical Simulations
Finally, we empirically study the value decomposition for the index policy on a circulant RMAB benchmark Avrachenkov and Borkar (2022), Zhang and Frazier (2022), Biswas et al. (2021), Fu et al. (2019) that has $4$ different states each local agent. As a result, the global state space scales as large as $4^{1800}>10^{1000}$ for $N=1800$ agents. The specific transitions and rewards are introduced in Appendix K. For each RMAB instance, we sample a trajectory of length $T=5N$ and use the collected data to i) solve Eq. (10) to estimate the measure of Markov entanglement; ii) train local Q-value decomposition. It quickly follows from the results in Figure 1:
@@ 7.2.1 Numerical Simulations
Figure 1: Circulant RMAB under an index policy. Left: empirical estimation of Markov entanglement multiplied by the number of agents, $NE_{1}({\bm{P}}^{\pi}_{1:N})$. Right: $\mu$-weighted decomposition error.
The estimated Markov entanglement decays as ${\mathcal{O}}(1/\sqrt{N})$ in the left panel, consistent with theoretical predictions. This also implies a low decomposition error scaling of ${\mathcal{O}}(\sqrt{N})$, as seen in the right panel. Furthermore, the simulated trajectory has a length of $T=5N$ while the global state space has size $|S|^{N}$, making both entanglement estimation and local Q-value decomposition sample-efficient.
@@ Appendix C Necessity of Negative Coefficients
We then construct the orthogonal complement of $\Omega_{P^{\otimes 2}}$ under Frobenius inner product. Let $\{\varepsilon_{j}\}_{j\in[m-1]}$ be a set of vector in $\mathbb{R}^{m}$ such that $\varepsilon_{j}=(1,0,\ldots,0,-1,0,\ldots,0)^{\top}$ with the first element $1$ and $j+1$-th element $-1$. Notice that
@@ Appendix E Decomposition via general functions
=$\stackrel{{\scriptstyle(i)}}{{=}}$ $\underbrace{({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}({\bm{r}}_{A}\otimes{\bm{e}})-\left({{\bm{I}}-\gamma{\bm{P}}_{A}\otimes{\bm{P}}_{B}}\right)^{-1}({\bm{r}}_{A}\otimes{\bm{e}})}_{(I)}$
@@ Appendix H Results for Multi-agent MDPs
Consider a $N$-agent MDP ${\mathcal{M}}_{1:N}$ and policy $\pi\colon{\mathcal{S}}\to\Delta({\mathcal{A}})$ with the measure of Markov entanglement $E_{i}(\pi)$ w.r.t ATV distance, it holds for any agent $i$,
$\left\lVert{\bm{P}}^{\pi}_{i}-{\bm{P}}_{i}\right\rVert_{\infty}\leq 2_{i}E(\pi)\,.$
@@
[Non-degenerate RMAB] There exists state $s\in{\mathcal{S}}$ such that $10<\pi^{\ast}(s,0)<$ 1, where $\pi^{\ast}$ is the policy under ${\bm{m}}^{\ast}$.
@@
Non-degenerate and UGAP are two standard technical assumptions for the index policy, which restrict chaotic behavior in asymptotic regime and will be further introduced in subsequent sections. We note here these two assumptions are also used in almost all theoretical work on index policies (Weber and Weiss 1990, Verloop 2016, Gast et al. 2023, 2024).
@@
Notice that the definition of $\phi^{\pi}$ does not depend on $N$. We adapt from Lemma B.1 in Gast et al. (2023) defined specially for Whittle Index,
@@
Lemma I.4 (One-step Concentration, Lemma 1 in Gast et al. (2024))
@@
Lemma I.5 (Multi-step Concentration, Lemma C.4 in Gast et al. (2023))
@@
Lemma I.6 (Local Stability, Lemma C.5 in Gast et al. (2023))
@@
=$\overset{(i)}{=}\Pr\left[{\left\lVert{\bm{m}}[\tilde{T}]-{\bm{m}}^{\ast}\right\rVert_{\infty}\geq{\epsilon}\mid{\bm{m}}[0]={\bm{m}}}\right]$
@@
In many practical scenarios, the agents’ transitions and actions are coupled by a shared exogenous signal. For example, in ride-hailing platforms, the specific dispatch is related to the exogenous order at the current moment (Qin et al. 2020, Han et al. 2022, Azagirre et al. 2024); in warehouse routing, the scheduling of robots is also related to the exogenous task revealed so far (Chan et al. 2024).
@@
Given the current state ${\bm{s}}$ and exogenous information $z$, the policy is given by $\pi:{\mathcal{S}}\times{\mathcal{Z}}\to\Delta(\tilde{\mathcal{A}})$, where $\tilde{\mathcal{A}}$ refers to the set of feasible actions. We then have the global transition depending on exogenous information $z$,
@@
$E(\pi,{\mathcal{Z}})\leq\frac{1}{2}\sup_{i}\sum_{s_{1:N},z}\mu^{\pi}(s_{1:N},z)\sum_{a_{i}}\left\lvert\pi(a_{i}\mid s_{1:N},z)-\pi^{\prime}(a_{i}\mid s_{i},z)\right\rvert\,,$
@@
Another common class of multi-agent MDPs is Factored MDPs (FMDPs, Guestrin et al. 2001, 2003, Osband and Roy 2014), which explicitly model the structured dependencies in state transitions. For instance, in a server cluster, the state transition of each server depends only on its neighboring servers. Formally, we define
@@
Although Proposition 1 offers a theoretical guarantee for general two-agent fully cooperative Markov games, its utility is greatest in systems with low reward and transition entanglement. Fully cooperative settings remain inherently challenging–for instance, even the asymptotically optimal Whittle Index may achieve only a $\frac{1}{N}$-approximation ratio for RMABs with global rewards (Raman et al. 2024). In practice, most research (Sunehag et al. 2018, Rashid et al. 2020) relies on sophisticated deep neural networks to learn decompositions in such settings. We thus defer a more refined analysis of fully cooperative scenarios to future work.
@@
In this section, we empirically study the value decomposition for index policies. Our simulations build on a circulant RMAB benchmark, which is widely used in the literature Avrachenkov and Borkar (2022), Zhang and Frazier (2022), Biswas et al. (2021), Fu et al. (2019).
@@
We set the discount factor to $\gamma=0.5$ and require $N/5$ arms to be pulled per period. Initially, there are $N/6$ arms in state 00, $N/3$ arms in state $1$ and $N/2$ arms in state $2$, the same as Zhang and Frazier (2022). We then test an index policy with priority: state $2>$ state $1>$ state $0>$ state $3$.
K.1 Monte-Carlo estimation of Markov entanglement
For each RMAB instance, we simulate a trajectory of length $T=6N$ and collect data for the later $5N$ epochs. Notice RMAB is a special instance of WCMDP, we thus apply the result in Proposition 7.2
$E_{i}({\bm{P}}_{1:N}^{\pi})$ $\leq\frac{1}{2}\min_{\pi^{\prime}}\sum_{{\bm{s}}}\mu^{\pi}_{1:N}({\bm{s}})\sum_{a_{i}}\left\lvert\pi(a_{i}\mid{\bm{s}})-\pi^{\prime}(a_{i}\mid s_{i})\right\rvert$
$\approx\frac{1}{2}\min_{\pi^{\prime}}\frac{1}{T}\sum_{t=1}^{T}\sum_{a_{i}}\left\lvert\pi(a_{i}\mid{\bm{s}})-\pi^{\prime}(a_{i}\mid s_{i})\right\rvert$ (15)
@@
K.2 Learning local Q-values
@@
The ground-truth $Q^{\pi}_{1:N}$ is approximated via Monte Carlo learning Sutton and Barto (2018), with each estimate derived from $30$-step simulations averaged over $3N$ independent runs. Due to the high computational cost of Monte Carlo methods—especially for very large RMABs—we limit the training phase to $10$ independent runs and use the mean local Q-value as an approximation. Error bars represent the standard error for both Monte Carlo estimates and $\mu$-weighted decomposition errors.
In addition to $\mu$-weighted error, we also introduce a concept of relative error, defined as $\left\lVert Q^{\pi}_{1:N}({\bm{s}},{\bm{a}})-\sum_{i=1}^{N}Q^{\pi}_{i}(s_{i},a_{i})\right\rVert_{\mu^{\pi}_{1:N}}/\left\lVert Q^{\pi}_{1:N}\right\rVert_{\mu^{\pi}_{1:N}}$. This relative error reflects the approximate ratio of our value decomposition. We present our simulation results in Figure 2.
@@
Figure 2: Value Decomposition error in circulant RMAB under an index policy. Left: $\mu$-weighted decomposition error. Right: Relative error, $\left\lVert\textrm{decomposition error}\right\rVert_{\mu}/\left\lVert Q^{\pi}_{1:N}\right\rVert_{\mu}$
It immediately follows that the $\mu$-weighted error grows at a sublinear rate ${\mathcal{O}}(\sqrt{N})$ and the relative error decays at rate ${\mathcal{O}}(1/\sqrt{N})$. This justifies our theoretical guarantees in Theorem 7.6. Furthermore, we notice the relative error is no larger than 3% over all data points. As a result, Meta Algorithm 1 is able to provide a very close approximation especially for large-scale MDPs even with small amount of training data $T=5N$ while the global state space has size $|S|^{N}$.
K.3 Sample Complexity and Computation
While each RMAB instance has an exponentially large state space $|S|^{N}$, we show that our empirical estimation of Markov entanglement—along with the decomposition error—converges quickly. Specifically, we illustrate these errors for an RMAB instance with with $900$ agents in Figure 3.

### v3

@@ 5.4 Distance Weighted by Stationary Distribution
In other words, the state action pair $({\bm{s}},{\bm{a}})$ with higher occupancy measure ${\mu^{\pi}_{AB}}({\bm{s}},{\bm{a}})$ contributes more to the overall decomposition error. We note that the error bound in $\mu$-norm is weaker than that in $\ell_{\infty}$-norm. Nevertheless, a stationary distribution weighted error bound is sufficient in many practical scenarios. Similar ideas are also quite common in policy evaluation literature (cai19neural, john96analysis, jalaj21afinite).
@@ 7 Applications of Markov Entanglement
This section discusses both theoretical and practical applications of Markov entanglement. We first show how specific MDP structures simplify entanglement analysis and produce sharp decomposition error bounds. We then demonstrate how Markov entanglement serves as an efficient test criteria of value decomposition for practitioners. Finally, we numerically study two important multi-agent scenarios: a synthetic restless multi-armed bandit model and a ride-hailing simulator.
@@
Weakly-coupled MDPs (WCMDP) are a rich class of multi-agent model that capture many real-world applications such as supply chain management, queuing network and resource allocations (daniel08relaxation, david23on, shar2023weakly). Compared to general multi-agent MDP, WCMDP further ensures each agent follow its local transition while the agents’ actions are coupled with each other. Formally,
@@
We demonstrate the proof for two-agent WCMDP and the generalization to multi-agent WCMDP is straightforward. Consider ${\bm{P}}_{A}^{\pi^{\prime}}$ be the transition of agent $A$ under local policy $\pi^{\prime}$. We focus on agent AA
@@
$\stackrel{{\scriptstyle(i)}}{{=}}\frac{1}{2}\sum_{{\bm{s}},{\bm{a}}}{\mu^{\pi}_{AB}}({\bm{s}},{\bm{a}})\sum_{s_{A}^{\prime},a_{A}^{\prime}}\left\lvert\sum_{s_{B}^{\prime}}P^{\pi}_{AB}({\bm{s}}^{\prime},a_{A}\mid{\bm{s}},{\bm{a}})-\sum_{s_{B}^{\prime}}P({\bm{s}}^{\prime}\mid{\bm{s}},{\bm{a}})\pi^{\prime}(a_{A}^{\prime}\mid s_{A}^{\prime})\right\rvert$
@@
$\stackrel{{\scriptstyle(ii)}}{{=}}\frac{1}{2}\sum_{{\bm{s}}^{\prime}}{\mu^{\pi}_{AB}}({\bm{s}}^{\prime})\sum_{a_{A}^{\prime}}\left\lvert\pi(a_{A}^{\prime}\mid{\bm{s}}^{\prime})-\pi^{\prime}(a_{A}^{\prime}\mid s_{A}^{\prime})\right\rvert\,$.
@@ 7.1.1 Index Policies are Asymptotically Separable
We further dive into the more structured multi-agent MDPs and introduce the following model of Restless Multi-Armed Bandit (RMAB), a special instance of weakly coupled MDP which is also widely used in operations research literature (Whittle88restless, weber90on, gast23exponential, zhang2021restless, zhang2022near).
@@ A Restless Multi-Armed Bandit is an $N$-agent WCMDP that further satisfies
  • There are two available actions for each agent: 0 for idle and $1$ for activate.
@@
In other words, RMAB is WCMDP with two actions and homogeneous agents that are coupled under constraint $\sum_{i=1}^{N}a_{i}=M$ at any global state $s_{1:N}$. This coupling of agents is often referred to budget constraint. For example, healthcare platform can reach out only a fraction of patients at a time due to the cost budget of interventions (baek2023policy).
@@
The index policies trace back to the well-known optimal Gittins Index (weber92on) and asymptotic optimal Whittle Index (Whittle88restless, weber90on, gast23exponential). More recent work generalizes Whittle Index to fluid based index policies (Verloop2016AsymptoticallyOP, gast24linear). However, computing the best index policy typically requires the knowledge of system transition a prior. As an alternative, qin20ride, xabi24better, baek2023policy, nakhleh21neu, wang2023optimistic, avrachenkov2022whittle apply data-driven method to optimize the index policy. Among these work, qin20ride, xabi24better, baek2023policy report great empirical success in industry-scale implementations. Understanding the mystery behind such success calls for a theory for general index policies. We then present our main result for this subsection
@@
Theorem 7.6 requires two standard technical conditions for index policies: non-degenerate and uniform global attractor property (UGAP), which restrict chaotic behavior in asymptotic regime and will be detailed in Appendix I. We note here these two assumptions are also used in many previous theoretical work on index policies (weber90on, Verloop2016AsymptoticallyOP, gast23exponential, gast24linear).
@@
This sublinear error result explains why the value decomposition in qin20ride, xabi24better, baek2023policy manages to effectively approximate the global value function in large-scale practical applications. It also justifies Meta Algorithm 1 as an effective approach to evaluating or comparing index policies.
@@
This lemma builds upon Proposition 7.2. We thus focus on the deviation from $\bm{m}$} to $\bm{m}$$^{\ast}$. We extend the concentration analysis from gast23exponential, gast24linear to derive a new stability bound for the RHS. Specifically, we finishing the proof via demonstrating the deviation decays at the rate ${\mathcal{O}}(1/\sqrt{N})$.
@@ 7.2 Efficient Verification of Value Decomposition
For practitioners, verifying the feasibility of value decomposition remains a significant challenge. Typically, they have to rely on indirect methods—for instance, evaluating policies derived from value decomposition in real-world settings or simulation environments, as seen in baek2023policy, qin20ride, xabi24better. While these policies often outperform baselines, it is unclear whether the learned decompositions accurately approximate the true global values. Moreover, real-world verifications are often prohibitively expensive. For example, the Lyft experiment (xabi24better) required a complex time-split design and posed substantial software engineering challenges in reliability and stability. Similarly, baek2023policy, qin20ride depended on extensive offline data collection to construct viable simulation environments.
@@
In other words, we can apply a Monte-Carlo estimation for estimating $E_{A}({\bm{P}}^{\pi}_{AB})$. Notice Eq. (9) is essentially a linear programming for ${\bm{P}}_{A}$ and this optimization can be solved distributively at each $(s_{A},a_{A})$ pair, which enables efficient solutions. Moreover, Eq. (9) only requires the knowledge of one-step transition ${\bm{P}}^{\pi}_{AB}(s_{A}^{\prime},a_{A}^{\prime}\mid{\bm{s}}^{t},{\bm{a}}^{t})$, which can often be easily calculated or simulated. For (weakly-)coupled MDPs, we can apply Proposition 7.2 and further eliminate the transition term,
@@
7.2.1 Numerical Simulation I: Restless Multi-armed Bandit
We first empirically study the value decomposition for the index policy on a circulant RMAB benchmark avrachenkov2022whittle, zhang2022near, biswas21learn, fu19towards that has $4$ different states each local agent. As a result, the global state space scales as large as $4^{1800}>10^{1000}$ for $N=1800$ agents. The specific transitions and rewards are introduced in Appendix K.1. For each RMAB instance, we sample a trajectory of length $T=5N$ and use the collected data to i) solve Eq. (10) to estimate the measure of Markov entanglement; ii) train local Q-value decomposition. It quickly follows from Figure 1:
@@ 7.2.1 Numerical Simulation I: Restless Multi-armed Bandit
Figure 1: Circulant RMAB under an index policy. Left: empirical estimation of Markov entanglement $E_{1}({\bm{P}}^{\pi}_{1:N})$. Right: $\mu$-weighted decomposition error.
The estimated Markov entanglement decays as ${\mathcal{O}}(1/\sqrt{N})$ in the left panel, consistent with theoretical predictions. This also implies a low decomposition error scaling of ${\mathcal{O}}(\sqrt{N})$, as seen in the right panel. Furthermore, the simulated trajectory has a length of $T=5N$ while the global state space has size $|S|^{N}$, showing both entanglement estimation and local Q-value decomposition sample-efficient.
7.2.2 Numerical Simulation II: A Ride-hailing Simulator
Finally, we study the Markov entanglement in ride-hailing (xabi24better, qin20ride), via a simulator built on NYC yellow cab data (NYCTLCData). The ride-hailing setting presents significantly greater complexity than RMAB. Most notably, a set of exogenous orders arrives at each timestep, and drivers are matched to these orders according to specific dispatching rules. Particularly, with $N$ drivers, we sample $0.1N$ orders at each timestep. Matched drivers transition to new positions based on their assigned orders, while idle drivers may also relocate autonomously (han22real). To address this challenge, we extend the Markov entanglement framework to accommodate exogenous orders (see Appendix J.1) and derive efficient estimators for both the Markov entanglement measure and local value functions in this setting. Simulation results are exhibited below with more details in Appendix K.2.
Refer to caption Refer to caption
Figure 2: A ride-hailing simulator. Left: empirical estimation of Markov entanglement $E_{1}({\bm{P}}^{\pi}_{1:N})$. Right: $\mu$-weighted decomposition error divided by the global value $\|Q^{\pi}_{1:N}\|_{\mu}$.
Perhaps surprisingly, despite the complexity of ride-hailing system, its Markov entanglement measure remains generally small and decays as ${\mathcal{O}}(1/\sqrt{N})$ for large $N$. We conjecture that this decay arises because the ride-hailing simulator converges to its mean-field limit as $N$ grows, which also exhibits asymptotic separability. More rigorous theoretical analysis is left for future work. Furthermore, the decomposition error relative to the true global Q-values decays at rate ${\mathcal{O}}(1/\sqrt{N})$, becoming negligibly small ($\leq 3\%$) for large $N$. These findings help explain the empirical success of value decomposition methods previously observed in ride-hailing applications.
@@ Appendix C Necessity of Negative Coefficients
We then construct the orthogonal complement of $\Omega_{P^{\otimes 2}}$ under Frobenius inner product. Let $\{\varepsilon_{j}\}_{j\in[m-1]}$ be a set of vector in $\mathbb{R}^{m}$ such that $\varepsilon_{j}=(1,0,\ldots,0,-1,0,\ldots,0)^{\top}$ with the first element $1$ and $j+1$-th element 1-1. Notice that
@@ Appendix E Decomposition via general functions
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\underbrace{({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}({\bm{r}}_{A}\otimes{\bm{e}})-\left({{\bm{I}}-\gamma{\bm{P}}_{A}\otimes{\bm{P}}_{B}}\right)^{-1}({\bm{r}}_{A}\otimes{\bm{e}})}_{(I)}$
@@ Appendix H Results for Multi-agent MDPs
Consider a $N$-agent MDP ${\mathcal{M}}_{1:N}$ and policy $\pi\colon{\mathcal{S}}\to\Delta({\mathcal{A}})$ with the measure of Markov entanglement $E_{i}({\bm{P}}^{\pi}_{1:N})$ w.r.t ATV distance, it holds for any agent $i$,
$\left\lVert{\bm{P}}^{\pi}_{i}-{\bm{P}}_{i}\right\rVert_{\infty}\leq 2_{i}E({\bm{P}}^{\pi}_{1:N})\,.$
@@
[Non-degenerate RMAB] There exists state $s\in{\mathcal{S}}$ such that $0<\pi^{\ast}(s,0)<1$, where $\pi^{\ast}$ is the policy under $\bm{m}$$^{\ast}$.
@@
Non-degenerate and UGAP are two standard technical assumptions for the index policy, which restrict chaotic behavior in asymptotic regime and will be further introduced in subsequent sections. We note here these two assumptions are also used in almost all theoretical work on index policies (weber90on, Verloop2016AsymptoticallyOP, gast23exponential, gast24linear).
@@
Notice that the definition of $\phi^{\pi}$ does not depend on $N$. We adapt from Lemma B.1 in gast23exponential defined specially for Whittle Index,
@@
Lemma I.4 (One-step Concentration, Lemma 1 in gast24linear)
@@
Lemma I.5 (Multi-step Concentration, Lemma C.4 in gast23exponential)
@@
Lemma I.6 (Local Stability, Lemma C.5 in gast23exponential)
@@
$\overset{(i)}{=}\Pr\left[{\left\lVert{\bm{m}}[\tilde{T}]-{\bm{m}}^{\ast}\right\rVert_{\infty}\geq{\epsilon}\mid{\bm{m}}[0]={\bm{m}}}\right]$
@@
In many practical scenarios, the agents’ transitions and actions are coupled by a shared exogenous signal. For example, in ride-hailing platforms, the specific dispatch is related to the exogenous order at the current moment (qin20ride, han22real, xabi24better); in warehouse routing, the scheduling of robots is also related to the exogenous task revealed so far (chan2024the).
@@
[…]
@@
$E_{i}({\bm{P}}^{\pi}_{1:N},{\mathcal{Z}})\leq\frac{1}{2}\sup_{i}\sum_{s_{1:N},z}\mu^{\pi}(s_{1:N},z)\sum_{a_{i}}\left\lvert\pi(a_{i}\mid s_{1:N},z)-\pi^{\prime}(a_{i}\mid s_{i},z)\right\rvert\,$, (14)
@@
Another common class of multi-agent MDPs is Factored MDPs (FMDPs, carlos01multiagent, carlos02efficient, ian14near), which explicitly model the structured dependencies in state transitions. For instance, in a server cluster, the state transition of each server depends only on its neighboring servers. Formally, we define
@@
Although Proposition J.7 offers a theoretical guarantee for general two-agent fully cooperative Markov games, its utility is greatest in systems with low reward and transition entanglement. Fully cooperative settings remain inherently challenging–for instance, even the asymptotically optimal Whittle Index may achieve only a $\frac{1}{N}$-approximation ratio for RMABs with global rewards (raman2024global). In practice, most research (peter18value, rashid2020monotonic) relies on sophisticated deep neural networks to learn decompositions in such settings. We thus defer a more refined analysis of fully cooperative scenarios to future work.
@@
All simulation code is available at this Github link.
K.1 Circulant RMAB
In this section, we empirically study the value decomposition for index policies.
@@
We set the discount factor to $\gamma=0.5$ and require $N/5$ arms to be pulled per period. Initially, there are $N/6$ arms in state 0, $N/3$ arms in state $1$ and $N/2$ arms in state $2$, the same as zhang2022near. We then test an index policy with priority: state $2>$ state $1>$ state $0>$ state $3$.
Monte-Carlo estimation of Markov entanglement
For each RMAB instance, we simulate a trajectory of length $T=6N$ and collect data for the later $5N$ epochs. Notice RMAB is a special instance of WCMDP, we thus apply Eq. (10).
$E_{i}({\bm{P}}_{1:N}^{\pi})\approx\frac{1}{2}\min_{\pi^{\prime}}\frac{1}{T}\sum_{t=1}^{T}\sum_{a_{i}}\left\lvert\pi(a_{i}\mid{\bm{s}})-\pi^{\prime}(a_{i}\mid s_{i})\right\rvert$ (16)
@@
Learning local Q-values
@@
The ground-truth $Q^{\pi}_{1:N}$ is approximated via Monte Carlo learning (Sutton18RL), with each estimate derived from $30$-step simulations averaged over $3N$ independent runs. Error bars represent the standard error for both Monte Carlo estimates and $\mu$-weighted decomposition errors.
In addition to $\mu$-weighted error, we introduce a concept of relative error, defined as $\left\lVert Q^{\pi}_{1:N}({\bm{s}},{\bm{a}})-\sum_{i=1}^{N}Q^{\pi}_{i}(s_{i},a_{i})\right\rVert_{\mu^{\pi}_{1:N}}/\left\lVert Q^{\pi}_{1:N}\right\rVert_{\mu^{\pi}_{1:N}}$. This relative error reflects the approximate ratio of our value decomposition. We present our simulation results in Figure 3.